# 100,000 Crab giant pulses at 215 MHz detected with an SKA-Low prototype station

### v1

@@ 6.1.3 Modelling scatter broadening time variations
Hence, the timescale of scattering the scattering is longer than the timescale of DM variations. Also, the scattering started and finished before the contribution of the plasma cloud to the total DM became measurable. It is expected that the onset of scattering becomes observable before the direct LoS is crossed by the cloud, and similar effect was observed by 2008A&A...483...13K.
We interpret our findings (Sections 6.1.2 and 6.1.3) using two screens model. With the variable component of the scattering caused by a screen in the expanding Crab Nebula or in the Perseus arm of the Galaxy, while the approximately constant “baseline” (2 ms) due to the second screen (not modelled here). Such a two screen model was suggested by the earlier studies (1971ApJ...166..513C; 1977ApJ...214..214I; 1975MNRAS.172...97L; 1976ApJ...209..578V; 1993ApJ...410..673G). The peaks (“horns”) in scattering time at around 43 – 58 and 65 – 83 days, may be due to additional sub-structures (lumps) of turbulent plasma within the main cloud. The fit to the data excluding these structures yielded very similar results. However, we defer more complex modelling using the two screen model, and potential inclusion of the “horns” to the future work.
@@ 6.1.4 Spectral behaviour of scatter broadening
We have also made an initial attempt to measure frequency scaling of scatter broadening time using bright GPs and software package SCAMP 999https://github.com/pulsarise/SCAMP-I (2021MNRAS.504.1115O; 2016MNRAS.462.2587G). The band was divided into 4 sub-bands and MCMC fitting of individual GP profiles was performed in each sub-band (see example in left Figure 13 for the brightest GP from 2024-12-24), and yielded scatter broadening times () consistent with our method. Based on this, scattering spectral index (where ) was fitted to scattering time as a function of frequency (e.g. right image in Figure 13). Figure 13 shows that the package performs well on Crab GPs despite the fact that it was originally designed to work on average pulse profiles. The same fitting procedure was applied to all the datasets and yielded as a function of time. The fitting performs very well during low scattering conditions for high SNR GPs, but did not converge well when the scattering was higher (SNR lower). Several outliers and failed fits were identified and discarded, and spectral index of the scatter time as a function of date is shown in Figure 14. The mean value of the spectral index is , which is in a good agreement with the previous measurements with the MWA (e.g. 2019ApJ...874..179K). Overall the spectral index varies significantly on various dates. Especially, the SCAMP fit did not perform well during high-scattering conditions regardless if isotropic or anisotropic scattering model was used. This could be caused by either the reduced SNR or neither of the models being applicable to these high scattering conditions. Additionally, the flattening of scattering spectral index observed when scattering was high (mid-February and early March 2025) may be a result of inadequate screen model.
These results show that any measurements of scattering spectral index for Crab pulsar have to be treated with caution. Mainly because highly variable scattering conditions may impact the fitting results on a particular day (e.g. cause apparent flattening) by reducing SNR or requiring different screen models (e.g. thin vs. thick, isotropic vs. anisotropic etc.). Additionally, as noted in 2021MNRAS.504.1115O; 2016MNRAS.462.2587G, the spectral index measurements are strongly dependent on the fitting procedure or package (e.g. fitting in log-log space vs. original values etc.). It is clear that consistent measurements require further work on screen models and more stable fitting procedures, especially during high-scattering/low-SNR conditions. We leave these efforts for the future, and we are willing to contribute to adjustments in the SCAMP package to make it suitable for single pulse analysis.
@@ 6.2.1 Spectral luminosity
The maximum spectral luminosities observed in our sample are of the order of ergs/s/Hz (right plot in Fig. 8), which is consistent with earlier measurements performed typically at higher frequencies (e.g. Figure 3 in 2022NatAs...6..393N). These spectral luminosities are some 3 orders of magnitude lower than the faintest pulses from repeating FRB 20200120E reported therein. By extrapolating the fitted spectral distribution of Crab GPs (right plot in Figure 8) to higher luminosities, we find that the waiting time for erg/s/Hz pulse is of the order of 66,000 years. The total number of galaxies up to redshift z8 was estimated by 2016ApJ...830...83C to be . We estimate number galaxies up to redshift z=0.1 to be (the highest reported FRB redshifts are 1 (2023Sci...382..294R)) by scaling by the ratio of co-moving volumes 101010https://www.astro.ucla.edu/~wright/CosmoCalc.html, while . We note that, scaling of by the ratio of co-moving volumes is a simplification not taking into account redshift evolution of the number of galaxies. Next, we assume number of young Crab-like pulsars per galaxy . There are no good estimates for this number and it could be as low as a 1 or as high as 100. Then a nearby FRB (z¡0.1) with spectral luminosity erg/Hz/s could be observed every few hours. Hence, FRBs similar to repeating FRB 20200120E, for which Crab-like pulsars are also considered as a potential progenitors 2021ApJ...910L..18B, could be relatively common. In fact, limits on number of pulses from such objects could provide estimates on the number of Crab-like pulsars.
Nevertheless, spectral luminosities of Crab GPs are very difficult to reconcile with some 15 orders higher spectral luminosities of typical one-off FRBs. Even assuming that the energy “reservoir” in Crab pulsar is sufficient to produce pulses ergs/s/Hz, the expected FRB wait times would about several orders of magnitude longer than the Hubble time. Hence, our observations are in agreement with the current models that Crab-like pulsars may be responsible for a population of weaker repeating FRBs, but cannot produce more energetic FRBs.
@@ 6.2.2 Arrival times
Several studies (2024ApJ...973...87D; 1995ApJ...453..433L; 2010A&A...515A..36K) showed that the distribution of pulse separation times for the Crab pulsar follows an exponential distribution, which corresponds to a purely random, Poisson, process. Hence, the detection of a pulse does not have impact on waiting time for the next one. This differs from what is observed in repeating FRBs, in which case the distribution is better characterised by Weibull distribution resulting from some form of correlation (clustering or anti-clustering) between the subsequent pulses (e.g. in clustering, detection of a pulse increases the probability of detecting another pulse). The distribution of time separations was studied before GPs (1995ApJ...453..433L; 2010A&A...515A..36K; 2024ApJ...973...87D). However, these earlier studies were performed at frequencies above 800 MHz (mainly L-band), and in some cases (e.g. 1995ApJ...453..433L) only up to a time difference of about 66 s (2000 pulsar periods). Therefore, we used our giant GP samples to test these hypotheses at low frequencies up to time separations of 600 s (vs. observation duration 3600 s).
@@
Figure 15 shows distributions of time between pulses for GPs with SNR10. We probed time differences up to 600 s calculated within each dataset only (up to 1 h observations), which ensured that the gaps between the observations or refractive scintillation did affect the analysis. The SNR10 threshold ensures that false-positives candidates, which can significantly impact the distribution, were not included. The resulting distribution can be reasonably well-fitted with an exponential distribution:
@@
resulting in (left image in Figure 15). The same fit was performed for each separate dataset, and in all cases the distribution of time between pulses was well described by an exponential distribution. The value of the fitted parameter varies between when the number of GPs detected per hour was highest (end of December 2024) down to 0.07 when the GP rate was lowest (middle February 2025). The plot of fitted vs. time has similar shape to Figure 4, and has not been included here for brevity. This varying behaviour of is perfectly in line with the fact that represents number of GPs in a certain time interval (GP rate) which was changing with changing scattering conditions (as discussed in Section 5.3).
@@
In order to look for potential deviations from a purely random process, the entire distribution was also fitted with a Weibull distribution (right image in Figure 15) :
which reduces to the exponential distribution for shape parameter k=1, and corresponds to clustering. The fit yielded k = , which significantly (18 standard deviations) differs from 1. The deviation from Poisson distribution (k=1) is relatively small, but statistically significant. This suggests that bright (SNR10) Crab GPs may exhibit some degree of clustering, and share some similarities with repeating FRBs (e.g. 2018MNRAS.475.5109O). Testing clustering of high-SNR GPs is justified by the fact that FRBs detected by radio-telescopes would correspond to the brightest GPs from extragalactic Crab-like pulsars.
@@ 6.2.3 Limits on low-frequency FRB rate
Similarly, no FRB-like events other than Crab giant pulses were detected in the entire dataset (repeating FRBs, RRATs, magnetars, Crab pulsar, nearby galaxy NGC 3109, and several other fields) totalling to about 100 hours of observations with the station beam FoV0.012 srad. This negative result can be translated into an upper limit on FRB rate Rfrb in 200 – 231.25 MHz band, which is R2500 FRBs/day/sky. This limit is some 2 orders of magnitude higher than the current extrapolations of FRB rates from higher frequencies (Figure 3 in 2024PASA...41...11S). Hence, the observing time and/or FoV need to increase substantially in order to routinely detect low-frequency FRBs. For example, even increasing number of beams to 10 could lead to first FRB detections after some 1000 hours of observations (about 6 weeks of non-stop observing with EDA2). However, millisecond all-sky imaging, as proposed by 2022aapr.confE...1S, is the best way to convert EDA2 and SKA-Low stations into “FRB machines” once the technical challenges of data capture, high-time resolution imaging, and real-time FRB searches are addressed (progress in these areas is described by 2024PASA...41...42S; 2025arXiv250515204D).
@@ 7 Summary and conclusions
Once more, standalone SKA-Low stations and EDA2 in particular turn out to be very flexible and productive scientific instruments for pulsar and transient sky monitoring. This work was possible thanks to the flexibility and quick turn around of the data as observations were typically processed and inspected within 24 hours from acquisition.
@@
During the observing campaign nearly 100,000 sample of giant pulses (GPs) from the pulsar B0531+21 (Crab) were detected, which is to our knowledge the largest GP sample at frequencies 300 MHz. In this paper, we presented statistical analysis of this unprecedented sample of GPs. Moreover, we see indications that time separation between high-SNR Crab GPs may be better characterised by Weibull (shape parameter k=) rather than exponential distribution. Hence, they may share some similarities with FRBs. These findings and the observed distribution of spectral luminosities are in-line with a possibility that extra-galactic Crab-like pulsars can produce weak repeating FRBs, but are unlikely progenitors of the entire population of FRBs. Given the observed distribution of spectral luminosities the wait times for typically bright FRBs from extra-galactic Crab-like pulsars would significantly exceed Hubble time. Hence, would be many orders of magnitude lower than the observed FRB rates (depending on the fluence, from 100s to 1000s per day per sky).
@@
We also observed a strong correlation of DM and scatter broadening time variations (correlation coefficient 0.7). Assuming that the event was caused by refractive scintillation, the estimated distance to plasma screen is of the order 1830 pc which places the screen in the Perseus arm of the Galaxy or even the Crab Nebula itself. The velocity of the scintillation pattern, km/s, is similar to the expansion velocities observed in the Crab Nebula, which further supports this possibility. Modelling of DM and variations leads to size of plasma screen pc with Gaussian distributed cloud of electron plasma of mean electron density e-cm-3, and peak electron density e-cm-3. This range of parameters is in agreement with earlier measurements, and matches electron densities in the filamentary structures of the Crab Nebula.
Our detections of Crab GPs (the faintest down to 500 Jy ms) clearly demonstrate that even a single SKA-Low station can be a powerful instrument in searching for and studying pulsars, FRBs and other fast transients. Especially, once multi-beaming capability is commissioned on the newly built SKA-Low stations (4 beams planned in Array Assembly 1 and more later). Furthermore, as discussed in 2024PASA...41...11S; 2022aapr.confE...1S; 2024PASA...41...42S the high time resolution (millisecond) imaging of the entire visible hemisphere, leveraged with the real-time FRB search pipelines (e.g. 2025arXiv250515204D) can yield even one FRB detection per day. The presented analysis demonstrates the first single-beam attempt and proof of principle for such efforts on a larger scale. Once multiple beams or thousands of image pixels with the same sensitivity can be formed, the scientific output from a single station can be boosted by orders of magnitude.
Our results provide a very optimistic prognostic for SKA-Low transient science, and of what may be possible in just a few years when hundreds of SKA-Low stations become operational. They also demonstrate SKA-Low potential to study local environments of pulsars, FRBs or other fast transients using low frequencies where propagation effects are much more pronounced than at higher frequencies. Given that nearby FRBs (z¡0.1) can be similarly bright, and their DMs similar or only a few times higher than Crab pulsar’s (e.g. FRB 20200120E has DM=87.82 pc cm-3 2021ApJ...910L..18B), we conclude that the huge potential of SKA-Low for wide-field and all-sky searches for FRBs and transients is only a few steps away!

### v2

@@ 6.1.3 Modelling scatter broadening time variations
Hence, the timescale of scattering the scattering is longer than the timescale of DM variations. Also, the scattering started and finished before the contribution of the plasma cloud to the total DM became measurable. It is expected that the onset of scattering becomes observable before the direct LoS is crossed by the cloud, and similar effect was observed by 2008A&A…483…13K (45).
We interpret our findings (Sections 6.1.2 and 6.1.3) using two screens model. With the variable component of the scattering caused by a screen in the expanding Crab Nebula or in the Perseus arm of the Galaxy, while the approximately constant “baseline” (2 ms) due to the second screen (not modelled here). Such a two screen model was suggested and applied by the earlier studies 1971ApJ…166..513C (18, 36, 56, 88, 31, 60). The peaks (“horns”) in scattering time at around 43 – 58 and 65 – 83 days, may be due to additional sub-structures (lumps) of turbulent plasma within the main cloud. The fit to the data excluding these structures yielded very similar results. However, we defer more complex modelling using the two screen model, and potential inclusion of the “horns” to the future work.
@@ 6.1.4 Spectral behaviour of scatter broadening
We have also made an initial attempt to measure frequency scaling of scatter broadening time using bright GPs and software package SCAMP 999https://github.com/pulsarise/SCAMP-I 2021MNRAS.504.1115O (70, 28). The band was divided into 4 sub-bands and MCMC fitting of individual GP profiles was performed in each sub-band (see example in left Figure 13 for the brightest GP from 2024-12-24), and yielded scatter broadening times () consistent with our method. Based on this, scattering spectral index (where ) was fitted to scattering time as a function of frequency (e.g. right image in Figure 13). Figure 13 shows that the package performs well on Crab GPs despite the fact that it was originally designed to work on average pulse profiles. The same fitting procedure was applied to all the datasets and yielded as a function of time. The fitting performs very well during low scattering conditions for high SNR GPs, but did not converge well when the scattering was higher (SNR lower). Several outliers and failed fits were identified and discarded, and spectral index of the scatter time as a function of date is shown in Figure 14. The mean value of the spectral index is , which is in a good agreement with the previous measurements with the MWA (e.g. 2019ApJ…874..179K (44)), and broadband measurements by 2006ARep…50..562P (72, 46). Overall the spectral index varies significantly on various dates. Especially, the SCAMP fit did not perform well during high-scattering conditions regardless if isotropic or anisotropic scattering model was used. This could be caused by either the reduced SNR or neither of the models being applicable to these high scattering conditions. Additionally, the flattening of scattering spectral index observed when scattering was high (mid-February and early March 2025) may be a result of inadequate screen model.
These results show that any measurements of scattering spectral index for Crab pulsar have to be treated with caution. Mainly because highly variable scattering conditions may impact the fitting results on a particular day (e.g. cause apparent flattening) by reducing SNR or requiring different screen models (e.g. thin vs. thick, isotropic vs. anisotropic etc.). Additionally, as noted in 2021MNRAS.504.1115O (70, 28), the spectral index measurements are strongly dependent on the fitting procedure or package (e.g. fitting in log-log space vs. original values etc.). It is clear that consistent measurements require further work on screen models and more stable fitting procedures, especially during high-scattering/low-SNR conditions. We leave these efforts for the future, and we are willing to contribute to adjustments in the SCAMP package to make it suitable for single pulse analysis.
@@ 6.2.1 Spectral luminosity
The maximum spectral luminosities observed in our sample are of the order of ergs/s/Hz (right plot in Fig. 8), which is consistent with earlier measurements performed typically at higher frequencies (e.g. Figure 3 in 2022NatAs…6..393N (66)). These spectral luminosities are some 3 orders of magnitude lower than the faintest pulses from repeating FRB 20200120E reported therein. By extrapolating the fitted spectral distribution of Crab GPs (right plot in Figure 8) to higher luminosities, we find that the waiting time for erg/s/Hz pulse is of the order of 66,000 years. The total number of galaxies up to redshift z8 was estimated by 2016ApJ…830…83C (14) to be . We estimate number galaxies up to redshift z=0.1 to be (the highest reported FRB redshifts are 1 2023Sci…382..294R (76)) by scaling by the ratio of co-moving volumes 101010https://www.astro.ucla.edu/~wright/CosmoCalc.html, while . We note that, scaling of by the ratio of co-moving volumes is a simplification not taking into account redshift evolution of the number of galaxies. Next, we assume number of young Crab-like pulsars per galaxy . There are no good estimates for this number and it could be as low as a 1 or as high as 100. Then a nearby FRB (z¡0.1) with spectral luminosity erg/Hz/s could be observed every few hours. Hence, FRBs similar to repeating FRB 20200120E, for which Crab-like pulsars are also considered as a potential progenitors 2021ApJ…910L..18B (6), could be relatively common. In fact, limits on number of pulses from such objects could provide estimates on the number of Crab-like pulsars.
Nevertheless, spectral luminosities of Crab GPs are very difficult to reconcile with some 15 orders higher spectral luminosities of typical one-off FRBs. Even assuming that the energy “reservoir” in Crab pulsar is sufficient to produce pulses ergs/s/Hz, the expected FRB wait times would be about several orders of magnitude longer than the Hubble time. Hence, our observations are in agreement with the current models that Crab-like pulsars may be responsible for a population of weaker repeating FRBs, but cannot produce more energetic FRBs.
@@ 6.2.2 Arrival times
Several studies 2024ApJ…973…87D (21, 54, 40) showed that the distribution of pulse separation times for the Crab pulsar follows an exponential distribution, which corresponds to a purely random, Poisson, process. Hence, the detection of a pulse does not have impact on waiting time for the next one. This differs from what is observed in repeating FRBs, in which case the distribution is better characterised by Weibull distribution resulting from some form of correlation (clustering or anti-clustering) between the subsequent pulses (e.g. in clustering, detection of a pulse increases the probability of detecting another pulse). The distribution of time separations of GPs was studied before 1995ApJ…453..433L (54, 40, 21). However, these earlier studies were performed at frequencies above 800 MHz (mainly L-band), and in some cases (e.g. 1995ApJ…453..433L (54)) only up to a time difference of about 66 s (2000 pulsar periods). Additionally, recent high-frequency (1.6 GHz) observations by 2023ApJ…959..111L (50) reported detections of multiple (up to 6) micro-bursts within a single pulse window (1 ms), largely exceeding numbers expected from a purely independent process. At our frequencies, however, micro-bursts within a pulse window (or even a few ms) are “blended” together due to large scatter broadening time. Hence, we used our giant sample of GPs to verify the distribution of pulse time separations at time scales longer than the pulsar period (verification of our merging procedure confirmed lack of pulses separated by less than 50 ms).
@@
Figure 15 shows distributions of time between pulses for GPs with SNR10, where we only used “good data sets” (at least 3000 s of non-flagged data with at least 2000 GPs detected) to probe time differences up to 600 s calculated within each dataset only (up to 1 h observations). This ensured that the gaps between the observations or refractive scintillation did affect the analysis. The SNR10 threshold ensures that false-positives candidates, which can significantly impact the distribution, were not included. However, we verified that the SNR thresholds higher than 7 give consistent results. The resulting distribution can be very well-fitted with an exponential distribution:
@@
resulting in (left image in Figure 15). The same fit was performed for each separate dataset, and in all cases the distribution of time between pulses was well described by an exponential distribution. The value of the fitted parameter varies between when the number of GPs detected per hour was highest (end of December 2024) down to 0.09 when the GP rate was lower (end of March 2025). The plot of fitted vs. time has similar shape to Figure 4, and has not been included here for brevity. This varying behaviour of is perfectly in line with the fact that represents number of GPs in a certain time interval (GP rate) which was changing with changing scattering conditions (as discussed in Section 5.3).
@@
The distribution was also fitted with a Weibull distribution (right image in Figure 15) :
which reduces to the exponential distribution for shape parameter k=1, and corresponds to clustering. The fit yielded k = consistent with 1, which confirms that the distribution of pulse arrival times is well fitted with exponential distribution as expected for independent events (i.e. Poisson process).
Finally, we have also created a distribution of logarithm to the base 10 of the wait times (Fig. 16), and compared the standard deviation of the fitted log-normal distribution to those expected for a periodic process, shot noise and observed in repeating FRBs (see Table 1 in 2024OJAp….7E..43K (41)). The standard deviation of the log-normal distribution fitted to our data is between 0 expected for a purely periodic process (e.g. regular pulses from a pulsar) and 0.723 expected for shot noise. However, it is lower than observed in repeating FRBs with sufficiently large number of pulses detected, and much lower than even observed in galactic magneters (see Table 1 in 2024OJAp….7E..43K (41)). This further confirms that statistical properties of arrival times of Crab GPs are different to those observed in repeating FRBs, which also indicates differences in physical emission mechanisms and processes.
@@ 6.2.3 Limits on low-frequency FRB rate
Similarly, no FRB-like events other than Crab giant pulses were detected in the entire dataset (repeating FRBs, RRATs, magnetars, Crab pulsar, nearby galaxy NGC 3109, and several other fields) totalling to about 100 hours of observations with the station beam FoV0.012 srad. This negative result can be translated into an upper limit on FRB rate Rfrb in 200 – 231.25 MHz band, which is R2500 FRBs/day/sky. This limit is some 2 orders of magnitude higher than the current extrapolations of FRB rates from higher frequencies (Figure 3 in 2024PASA…41…11S (78)). Hence, the observing time and/or FoV need to increase substantially in order to routinely detect low-frequency FRBs. For example, even increasing number of beams to 10 could lead to first FRB detections after some 1000 hours of observations (about 6 weeks of non-stop observing with EDA2). However, millisecond all-sky imaging, as proposed by 2022aapr.confE…1S (83), is the best way to convert EDA2 and SKA-Low stations into “FRB machines” once the technical challenges of data capture, high-time resolution imaging, and real-time FRB searches are addressed (progress in these areas is described by 2024PASA…41…42S (81, 20)).
@@ 7 Summary and conclusions
Once more, standalone SKA-Low stations and EDA2 in particular turned out to be very flexible and productive scientific instruments for pulsar and transient sky monitoring. This work was possible thanks to the flexibility and quick turn around of the data as observations were typically processed and inspected within 24 hours from acquisition.
@@
During the observing campaign nearly 100,000 sample of giant pulses (GPs) from the pulsar B0531+21 (Crab) were detected, which is to our knowledge the largest GP sample at frequencies 300 MHz. In this paper, we presented statistical analysis of this unprecedented sample of GPs. On the timescales accessible to our analysis (1 ms), the observed distribution of pulse arrival times is consistent with exponential distribution expected for purely random, Poisson process. These findings are consistent with the earlier observations that Crab GPs, unlike FRBs, do not show any signs of clustering observed in FRBs. Due to limitations of low-frequency observations, further high time resolution, high-frequency studies will be required to confirm the clustering of micro-bursts reported by 2023ApJ…959..111L (50). Our findings and the observed distribution of spectral luminosities do not exclude a possibility that extra-galactic Crab-like pulsars can produce weak repeating FRBs, but are unlikely progenitors of the entire population of FRBs. Especially, that given the observed distribution of spectral luminosities the wait times for typically bright FRBs from extra-galactic Crab-like pulsars would significantly exceed Hubble time. Hence, the FRB rates would be many orders of magnitude lower than the observed rates (depending on the fluence, from 100s to 1000s per day per sky).
@@
We also observed indications of correlation of DM and scatter broadening time variations (correlation coefficient 0.7), which may be overestimated due to the impact of scatter broadening on DM measurements.. Assuming that the event was caused by refractive scintillation, the estimated distance to plasma screen is of the order 1830 pc which places the screen in the Perseus arm of the Galaxy or even the Crab Nebula itself. The velocity of the scintillation pattern, km/s, is similar to the expansion velocities observed in the Crab Nebula, which further supports this possibility. Modelling of DM and variations leads to size of plasma screen pc with Gaussian distributed cloud of electron plasma of mean electron density e-cm-3, and peak electron density e-cm-3. This range of parameters is in agreement with earlier measurements, and matches electron densities in the filamentary structures of the Crab Nebula.
Our detections of Crab GPs (the faintest down to 500 Jy ms) clearly demonstrate that even a single SKA-Low station can be a powerful instrument in searching for and studying pulsars, FRBs and other fast transients. Especially, once multi-beaming capability is commissioned on the newly built SKA-Low stations (4 beams planned in Array Assembly 1 and more later). Furthermore, as discussed in 2024PASA…41…11S (78, 83, 81) the high time resolution (millisecond) imaging of the entire visible hemisphere, leveraged with the real-time FRB search pipelines (e.g. 2025arXiv250515204D (20)) can yield even one FRB detection per day. The presented analysis demonstrates the first single-beam attempt and proof of principle for such efforts on a larger scale. Once multiple beams or thousands of image pixels with the same sensitivity can be formed, the scientific output from a single station can be boosted by orders of magnitude.
Our results provide a very optimistic prognostic for SKA-Low transient science, and of what may be possible in just a few years when hundreds of SKA-Low stations become operational. They also demonstrate SKA-Low potential to study local environments of pulsars, FRBs or other fast transients using low frequencies where propagation effects are much more pronounced than at higher frequencies. Given that nearby FRBs (z¡0.1) can be similarly bright, and their DMs similar or only a few times higher than Crab pulsar’s (e.g. FRB 20200120E has DM=87.82 pc cm-3 2021ApJ…910L..18B (6)), we conclude that the huge potential of SKA-Low for wide-field and all-sky searches for FRBs and transients is only a few steps away!